Determine a region of the xy-plane for which the given differential equation would have a unique solution whose graph passes through a point (x0, y0) in the region.

(25 − y2)y' = x2

Choose the right answer and explain

a. A unique solution exists in the regions y < −5, −5 < y < 5, and y > 5.

b. A unique solution exists in the region y < 5.

c. A unique solution exists in the region consisting of all points in the xy-plane except (0, 5) and (0, −5).

d. A unique solution exists in the region y > −5.

e. A unique solution exists in the entire xy-plane.
Answer #1

23 In the (25-ya) y = x² ya f (2) y (6) - Yo (25-y2) IVP y = f (x, y, y (ro) - Yo Is fray) and continuous function in some o